# SpeechRole: A Large-Scale Dataset and Benchmark for Evaluating Speech Role-Playing Agents

### v4

@@ 4.2 Evaluation Metrics
In contrast to traditional text-based RPAs, SRPAs must not only accurately emulate specific characters but also consistently deliver coherent and expressive spoken responses fully aligned with both the dialogue context and the character’s unique persona. To comprehensively evaluate SRPAs, we propose a robust three-dimensional evaluation framework that assesses fundamental interaction ability, speech quality and expressiveness, and role-playing fidelity.
Fundamental interaction abilities comprehensively evaluate the agent’s capacity to maintain coherent, natural, and instruction-aligned interactions fully consistent with the character, which are assessed by the following key metrics:
@@
Speech quality and expressiveness measure the naturalness and prosodic expressiveness of the generated speech, which are assessed by the following metrics:
@@
Role-playing fidelity assesses the extent to which the agent embodies the assigned character’s persona, which is assessed by the following metrics:
@@
4.3 Evaluation Procedure
LLM-Based Scoring Procedure.
There is a lack of widely recognized objective metrics for the evaluation of role speech output, and direct scoring by LLMs often exhibits instability and inconsistency [DBLP:conf/acl/WangLCCZLCKLLS24]. Inspired by DBLP:conf/iclr/DaiH0J0025, two strategies are adopted to enhance the reliability of LLM-based scoring.
First, the model is required to generate a brief rationale for its score prior to providing the original rating. This process encourages more deliberate and interpretable evaluations by the model. Second, rather than evaluating individual speech samples in isolation, a paired comparison setting is designed. For each evaluation instance, the model receives a test speech sample and a high-quality reference answer simultaneously, and assigns scores to both.
The final evaluation score is calculated as the ratio of the test speech score to the reference score, which standardizes the results across different prompts and reduces variance caused by prompt discrepancies. This two-step, reference-based evaluation framework is designed to improve scoring stability and promote better alignment with human judgments. We use gemini-2.5-pro-preview-06-05 [DBLP:journals/corr/abs-2312-11805] to assign scores to each response according to the predetermined evaluation protocol, with scores ranging from 1 to 10.
Human Correlation Verification.
To verify the reliability of our automatic scoring method, we randomly select 20 dialogue samples and, for each sample, randomly choose two model-generated responses. Four independent human annotators conduct pairwise comparisons of these responses based on multiple evaluation metrics, determining which response performs better for each metric. The final human score for each metric is calculated as the average of all annotators’ judgments. The results show a strong correlation between human evaluation and automatic scoring, confirming that our scoring method aligns closely with human judgments and provides a reliable foundation for large-scale automated evaluation. The human judgment setting follows the evaluation protocol introduced by DBLP:conf/iclr/DaiH0J0025, which uses pairwise preference comparisons to ensure stable and interpretable annotations. Details are in the appendix.
Models Evaluation Metrics Overall
IA SF $C$ SN PC EA PeC KC
English Evaluation Results
Alibaba Cloud API 0.9012 0.6678 1.0320 0.6973 0.6651 0.8087 0.8594 0.9171 0.8186
XTTS 0.9617 0.5493 0.9872 0.6578 0.7568 1.0133 1.0357 0.9531 0.8644
CosyVoice2 1.0102 0.7869 1.0346 0.8682 0.9857 1.0781 1.0476 0.9744 0.9732
E2 TTS 1.0531 0.7938 1.1621 0.8859 1.0718 1.1806 1.1760 1.0104 1.0417
F5-TTS 1.1245 0.8057 1.1132 0.8717 0.9773 1.2686 1.0952 1.0614 1.0397
FireRedTTS 0.9587 0.7266 1.0297 0.8135 0.8760 1.0557 1.0415 0.9899 0.9365
Spark-TTS 1.0384 0.7583 1.0641 0.8282 1.0554 1.1427 1.0741 1.0144 0.9970
GPT-SoVITS 0.8566 0.4783 0.9293 0.5576 0.5606 0.8253 0.8899 0.8770 0.7468
MaskGCT 0.9277 0.7635 0.9902 0.7589 0.9449 1.0382 0.9701 1.0098 0.9254
[…]
Chinese Evaluation Results
Alibaba Cloud API 0.8988 0.8473 0.8942 0.9275 1.0223 1.0493 0.9124 0.8887 0.9301
XTTS 0.6563 0.4951 0.7149 0.5348 0.4681 0.4945 0.6191 0.6595 0.5803
CosyVoice2 0.7600 0.7739 0.7190 0.7943 0.7388 0.8843 0.6587 0.6597 0.7486
E2 TTS 0.6866 0.6513 0.7171 0.7082 0.6776 0.7027 0.6787 0.7161 0.6923
F5-TTS 0.7285 0.6901 0.7658 0.8052 0.8392 0.9291 0.6969 0.6890 0.7680
FireRedTTS 0.7536 0.6932 0.6968 0.7221 0.8371 0.8235 0.7093 0.6516 0.7359
Spark-TTS 0.7591 0.7179 0.7869 0.7887 0.8930 0.7902 0.7268 0.6890 0.7690
GPT-SoVITS 0.6915 0.6345 0.7269 0.6119 0.6583 0.6515 0.6296 0.6277 0.6540
MaskGCT 0.7391 0.5913 0.6558 0.7349 0.8011 0.8109 0.7722 0.6433 0.7186
[…]
Table 2: Comprehensive evaluation results of $1$ SRPAs on the SpeechRole-Eval benchmark. The table reports model performance across eight evaluation metrics: IA (Instruction Adherence), SF (Speech Fluency), $C$ (Conversational Coherence), SN (Speech Naturalness), PC (Prosody Consistency), EA (Emotion Appropriateness), PeC (Personality Consistency), and KC (Knowledge Consistency), along with the overall average score. Results are presented for both English and Chinese test sets. End-to-end models are indicated with a gray background. Proprietary models include Alibaba Cloud API and GPT-4o Audio, while all others are open-source.
5 Experiments
This section details our experimental setup and main results. We evaluate SRPAs under two paradigms: cascaded systems, which separate role-playing into ASR, LLM-based reasoning, and TTS stages, and end-to-end models that generate speech directly from prompts. All models are tested on the SpeechRole-Eval benchmark (Section 4) covering eight core metrics, with speaker similarity assessed for voice-controllable models. We also include supplementary experiments, such as fine-tuning a SRPA on our dataset to assess its speech role-playing performance.
5.1 SRPA Architectures
The primary objective of SRPAs is to generate coherent, spoken responses that are consistent with the target character’s identity upon receiving a user’s spoken query. To explore this, we examine two representative system architectures: the cascaded and end-to-end approaches.
Cascaded SRPAs decompose the overall task into three modular components. First, an ASR model transcribes the user’s spoken input into text. Next, a LLM generates a contextually appropriate and character-specific text-based response that aligns with the specified character’s identity, using both the transcription and system prompts defining the role. Finally, a TTS module converts the generated text into speech that matches the target character’s voice. In our implementation, both proprietary and open-source cascaded systems are explored. We utilize a proprietary cascaded system based on the Alibaba Cloud API [alibabacloud], integrating Paraformer Realtime ASR v2, Qwen-Plus-Character, and CosyVoice-v2. For the open-source side, we employ whisper-large-v3-turbo [DBLP:conf/icml/RadfordKXBMS23] for high-quality ASR, LLaMA3.1-8B [DBLP:journals/corr/abs-2407-21783] for flexible and powerful text generation, and multiple state-of-the-art zero-shot timbre cloning TTS models for natural and expressive speech synthesis, including XTTS [DBLP:conf/interspeech/CasanovaDGGGHAM24], CosyVoice2 [DBLP:journals/corr/abs-2412-10117], E2 TTS [DBLP:conf/slt/EskimezWTLTXYZTTLZK24], F5-TTS [DBLP:journals/corr/abs-2410-06885], FireRedTTS [DBLP:journals/corr/abs-2409-03283], GPT-SoVITS [gpt-sovits], Spark-TTS [DBLP:journals/corr/abs-2503-01710], and MaskGCT [DBLP:conf/iclr/WangZLZGZZZZ025].
End-to-end SRPAs generate speech responses directly from spoken input, guided by the user’s spoken query and a textual prompt of the character’s identity. Both the proprietary gpt-4o-audio-preview-2024-12-17 [DBLP:journals/corr/abs-2410-21276] and the open-source Qwen2.5-Omni-7B [DBLP:journals/corr/abs-2503-20215] are evaluated in our study.
5.2 Evaluation Results
Main Evaluation Results.
We report the main evaluation results of all SRPAs on the SpeechRole-Eval benchmark. This benchmark covers eight core metrics: Instruction Adherence (IA), Speech Fluency (SF), Conversational Coherence (CC), Speech Naturalness (SN), Prosody Consistency (PC), Emotion Appropriateness (EA), Personality Consistency (PeC), and Knowledge Consistency (KC), with the final score computed as the average across these dimensions.
As shown in Table 2, the cascaded agents generally outperform end-to-end systems across most evaluation dimensions. Notably, models such as E2 TTS, F5-TTS, and Spark-TTS achieve relatively high overall scores in both content relevance and expressiveness, benefiting greatly from their modular design and optimization. While end-to-end models, particularly GPT-4o Audio, demonstrate competitive performance in expressiveness and coherence, they still exhibit noticeable shortcomings in style consistency and character setting. The performance of Qwen2.5-Omni is significantly lower across all metrics, indicating that end-to-end modeling remains challenging in the absence of sufficient task-specific supervision.
Language-Level Observations.
We observe significant differences between the English and Chinese subsets. Overall, the model performs better on English, particularly with respect to metrics such as content coherence and knowledge consistency. This disparity likely reflects the model’s language imbalance in pre-training and TTS capabilities. For example, XTTS performs relatively well in English but shows a marked decline in Chinese, especially in the expression of emotion and character personality. In contrast, Alibaba Cloud API and CosyVoice2 demonstrate more stable performance across both languages. These results indicate that current SRPAs exhibit uneven performance across different languages, with significantly better results in English than in Chinese. Therefore, further efforts are necessary to develop high-quality multilingual SRPAs.
Figure 4: Speaker similarity scores of SRPAs with voice cloning capability.
Speaker Similarity Comparison.
We further analyze the performance of SRPAs with voice cloning capability in terms of speaker similarity. Since end-to-end models do not support speaker cloning, we evaluate speaker similarity only for cascade systems. Following the same evaluation protocol as Seed-TTS [DBLP:journals/corr/abs-2406-02430], we use the WavLM-large [DBLP:journals/jstsp/ChenWCWLCLKYXWZ22] model to extract speaker embeddings and compute cosine similarity scores.
As shown in Figure 4, in the English setting, E2 TTS and F5-TTS achieve the highest speaker similarity, while GPT-SoVITS performs the worst. The overall ranking of models remains largely consistent across the two languages, indicating stable voice cloning performance. Notably, MaskGCT slightly outperforms others in Chinese, reaching the highest SIM score.
Models Evaluation Metrics Overall
IA SF $C$ SN PC EA PeC KC
English Evaluation Results
Qwen2.5-Omni 0.3977 0.4968 0.5491 0.4505 0.2958 0.3814 0.4089 0.4917 0.4340
SpeechRole-Agent 0.8072 0.6790 0.9480 0.6123 0.4202 0.5383 0.7568 0.9169 0.7098
[…]
Chinese Evaluation Results
Qwen2.5-Omni 0.4924 0.6886 0.5832 0.6204 0.4693 0.5125 0.3932 0.4772 0.5296
SpeechRole-Agent 0.9188 0.8233 0.9037 0.7749 0.7277 0.8607 0.8402 0.9253 0.8468
[…]
5.3 Fine-Tuning SRPA
Training Setup.
We fine-tune Qwen2.5-Omni-7B on the SpeechRole training set, containing 78 roles with speech-based dialogues. The model trains using speech inputs paired with role-consistent responses, enabling it to better understand and generate role-based spoken interactions. During evaluation, we use SpeechRole-Eval, dividing it into two subsets: in-domain, which includes roles seen during training, and out-of-domain, consisting of unseen roles. Results are reported separately for English and Chinese.
Performance Analysis.
After fine-tuning on the complete SpeechRole-Data training set, SpeechRole-Agent achieves notable gains over the original Qwen2.5-Omni. In the in-domain test set, the overall score in English increases from 0.4340 to 0.7098, while in Chinese it rises from 0.5296 to 0.8468. All evaluation metrics improve, with particularly notable advancements in instruction adherence, emotional appropriateness, and personality consistency. In the out-of-domain test set, the model also demonstrates strong generalization ability, with the English score increasing from 0.4786 to 0.6795 and the Chinese score from 0.5117 to 0.8028. These results fully demonstrate the model’s robust role-playing capabilities beyond the scope of the training set. Detailed results are presented in Table 3.
6 Conclusion
We propose SpeechRole, a unified framework for developing and evaluating SRPAs. To support this objective, we construct SpeechRole-Data, a large-scale, high-quality dataset comprising 112k dialogues from 98 characters, each with rich role settings and diverse vocal characteristics. Our benchmark establishes rigorous evaluation protocols for role-consistent speech generation and introduces SpeechRole-Eval, an automated multimodal evaluation suite that assesses various aspects including interaction competence, speech expressiveness, and role consistency.
Experimental results indicate current SRPAs still face challenges in achieving fully coherent and expressive role simulation, underscoring the need for more advanced character-based speech behavior modeling. We expect this work will provide a solid foundation for future research at the intersection of speech generation, role-playing, and embodied agents. All data, evaluation protocols, and baseline methods will be made publicly available to facilitate further studies on speech-interactive agents.
Figure 5: Left: Distribution of eight sub-tasks across three major task categories in SpeechRole. Right: Distribution of the 98 roles in SpeechRole-Data by language, gender, and data source.
Figure 6: Main evaluation results of $1$ SRPAs on SpeechRole-Eval. Left: English results. Right: Chinese results.Each metric is displayed with an interval range of 0.26, with all evaluation results normalized such that the maximum value across metrics is scaled to 1.30.
Appendix A SpeechRole-Data Construction
A.1 Text Data Construction
Role Selection.
A total of 98 representative roles are meticulously curated for SpeechRole-Data, selected from two high-quality datasets: ChatHaruhi [DBLP:journals/corr/abs-2308-09597] and RoleLLM [DBLP:conf/acl/WangPQLZWGGN00024]. Specifically, 18 roles from games, movies, and television series are chosen from ChatHaruhi, while 80 roles from animation, movies, and television series are drawn from RoleLLM. The selected roles are balanced in terms of language, gender, and personality traits, thereby ensuring diversity in both vocal expression and character identity (see Figure 5). To facilitate rich and character-consistent speech modeling, particular emphasis is placed on roles with distinctive vocal characteristics. Table 5 and Table 6 summarizes the attributes of all roles, including their categories, languages, genders, source materials, and dataset splits.
Role Metadata Extraction.
For each character, we extract structured metadata, including: (1) Role Profile, such as temperament, behavioral tendencies, values, and personal preferences; (2) Background, including character settings, social identity, relationships, and living environment; and (3) Character Lines, including narration, monologues, and dialogue excerpts involving the target character in games or audiovisual works. These elements collectively define distinctive character profiles and support vivid, context-rich speech role-playing. Additionally, the dialogue excerpts serve as the basis for guiding dialogue generation and evaluating character consistency.
Dialogue Generation.
We use gpt-4.1-2025-04-14 [DBLP:journals/corr/abs-2303-08774] to generate dialogues between users and roles based on role metadata. For some roles with lengthy scripts, the content is divided into segments and processed sequentially. GPT-4.1 is required to provide relevant script segments while generating dialogues to serve as references for SRPA in answering questions. For each role, 800 single-turn dialogues and 800 multi-turn dialogue data are generated.
Dialogue Filtering.
The data are filtered through a multi-stage process to ensure structural consistency, contextual relevance, and semantic diversity. First, only dialogues that conform to a clear turn-taking pattern are retained. In these dialogues, the user initiates with a question, and the role responds, with each party speaking in alternating turns. To ensure alignment between the dialogue and its corresponding context, semantic embeddings are computed for both texts. For English data, the all_MiniLM_L6_v2 [DBLP:conf/emnlp/ReimersG19] model is used, and samples with a similarity score below 0.4 are excluded. The same procedure is applied to Chinese data using the text2vec-bge-large-chinese [text2vec] model, with a similarity score threshold of 0.45.
To further eliminate semantic redundancy, which is commonly observed in GPT-4.1-generated dialogues, near-duplicate samples are removed based on the similarity of their dialogue embeddings. Thresholds are set to 0.85 for Chinese and 0.9 for English. This de-duplication step is conducted separately for single-turn and multi-turn dialogues for each role. After filtering, the dataset contains 53,902 single-turn dialogues and 57,944 multi-turn dialogues.
Metrics IA SF $C$ SN PC EA PeC KC Overall
MSE 0.1020 0.0413 0.1426 0.0581 0.2539 0.1262 0.1690 0.1147 0.1260
Pearson 0.7304 0.6366 0.6076 0.8093 0.6316 0.8881 0.6677 0.7581 0.7162
Table 4: Evaluation results of correlation between automatic and human scoring.
A.2 Role Voice Collection
Role Voice Extraction.
To achieve authentic voice role-playing, real-life performance audio for each character is collected as reference material. For characters from Genshin Impact, in-game character audio is collected with permission from miHoYo. For other characters, audio is obtained from actual film and television works. Films and television series featuring these characters are gathered, and a total of 105 video files with an approximate duration of 181 hours are downloaded from the internet, covering all target characters.
During the preprocessing stage, all extracted speech segments undergo anonymization and fragmentation to ensure that they contain only brief, decontextualized utterances, with no retention of the original video or narrative context. As a result, no identifiable or copyrighted content remains, thus mitigating potential intellectual property risks and ensuring adherence to the principles of fair use for research purposes.
We first employ a batch processing pipeline based on ffmpeg [ffmpeg] to extract the raw audio streams from the collected video files. To ensure consistency, all audio is converted to the standard mono 16 kHz WAV format. However, since these recordings often contain background music, noise, and overlapping speech, additional processing is required.
Role Voice Processing.
To obtain clean audio segments from specific speakers, we employ a structured pipeline based on the open-source Emilia [DBLP:conf/slt/HeSWLGHLYLSWCZW24] framework. This process includes source separation to remove background noise, speaker separation to distinguish individual speakers, and voice activity detection for fine-grained segmentation. The resulting audio segments, each lasting between 3 and 10 seconds and containing only a single speaker, are well-suited for subsequent modeling.
Finally, we use an automatic speech recognition model to transcribe each segmented audio fragment and evaluate speech quality using the DNSMOS P.835 OVRL [DBLP:conf/icassp/ReddyGC22] metric. In order to ensure the quality of speech data, we filter out speech with DNMOS less than 3. The resulting structured dataset provides, for each audio entry, the following information: start and end times, transcription, speaker index, language, and DNSMOS-based audio quality score.
Speaker Identification.
The extracted audio data contains only the speaker’s index and does not include identity information. To identify the speaker, we use both gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [DBLP:journals/corr/abs-2412-19437] models to determine the target character’s speaker number based on all transcribed texts. We integrate the results from both models and verify them manually to obtain the speech corresponding to the target character.
Role Voice Ranking.
To select the most representative audio sample as a reference for subsequent role-playing, we rank all speech samples of each speaker based on similarity. Specifically, we extract the speaker feature vector of each audio sample using the CAM++ [DBLP:conf/interspeech/WangZCC023] model and calculate the cosine similarity between each sample and all other samples. A higher average similarity score indicates that the audio sample better represents the overall speaker characteristics of the role. Finally, we sort all audio samples for each role in descending order according to their similarity scores and select the top-ranked sample as the reference audio.
Appendix B Human Correlation Verification Details
To evaluate the consistency between automatic scoring and human scoring, we randomly selected 20 samples from the dataset, including 5 single-turn dialogue samples and 15 multi-turn dialogue samples. For each sample, we randomly selected two model-generated speech responses for evaluation.
Gemini scored each model response and its corresponding reference response on a scale of 1 to 10 based on 8 predefined scoring metrics. The final score for a model was calculated as the ratio of the model response score to the reference response score.
For human evaluation, four trained evaluators were employed to directly compare the responses of the two models based on the same 8 scoring metrics. The evaluation criteria were as follows: the superior model received 1 point, the inferior model 0 points, and if the performance was deemed equivalent, each received 0.5 points. The final human score was determined by averaging the scores from all evaluators.
To quantify the agreement between automated and human evaluation methods, we computed two key metrics: mean squared error (MSE) and Pearson correlation coefficient. The MSE was calculated by first fitting a linear regression between the score differences of paired model responses from human evaluators and gemini. The predicted gemini score differences were then derived from this regression model, and the MSE was computed between these predicted values and the actual gemini score differences. For the Pearson correlation analysis, we computed the linear correlation coefficient between the human evaluators’ score differences and gemini’s score differences across all paired comparisons.
Table 4 reports the detailed results, the overall MSE of 0.1260 indicates relatively minor deviations between the two scoring methods. The high Pearson correlation coefficient further confirms the reliability of the automated scoring system. The evaluation results collectively demonstrate strong agreement between automatic and human scoring across all metrics, indicating that the automated scoring system can effectively approximate human judgment.

### v6

@@ 4.1 Evaluation Metrics
SRPAs must handle not only the linguistic demands of dialogue generation but also the vocal and stylistic requirements of spoken character portrayal. Evaluating such systems therefore requires attention to aspects beyond those considered in traditional text-based settings. To capture the full range of SRPA capabilities, we adopt a three-dimensional evaluation framework.
Interaction Ability: measures the agent’s capacity to produce coherent, contextually appropriate, and instruction-aligned responses. Assessed using Instruction Adherence (IA) and Conversational Coherence (CC).
Speech Quality and Expressiveness: evaluates the naturalness, fluency, and prosodic expressiveness of speech. Assessed using Speech Fluency (SF), Speech Naturalness (SN), Prosodic Consistency (PC), and Emotion Appropriateness (EA).
Role-Playing Fidelity: examines how well the agent maintains character-specific personality and knowledge. Assessed using Personality Consistency (PeC) and Knowledge Consistency (KC).
4.2 LLM-Based Evaluation Procedure
There is a lack of widely recognized objective metrics for evaluating role-conditioned speech outputs, and absolute scoring by LLMs is known to be unstable and inconsistent, largely due to the open-ended nature of role-playing, the absence of a single correct response, and the sensitivity of absolute scores to prompt phrasing and scoring scale [DBLP:conf/acl/WangLCCZLCKLLS24]. To address these limitations, recent work advocates pairwise scoring as a more reliable alternative [DBLP:conf/iclr/DaiH0J0025], as it bypasses the need for a ground-truth answer, improves inter-judge consistency, and produces preference signals that better align with human judgments. Motivated by these findings, SpeechRole-Eval adopts a pairwise relative scoring protocol using an LLM with strong speech understanding capabilities, enabling stable, objective, and reproducible evaluation of SRPAs.
To enhance evaluation consistency and interpretability, the LLM is first prompted to produce a concise rationale before issuing its rating, encouraging more deliberate judgment. Moreover, instead of assessing each speech sample independently, we adopt a pairwise comparative setting, where the LLM jointly receives a test sample and a high-quality reference response and assigns scores to both, enabling more stable and discriminative relative assessments.
We employ gemini-2.5-pro [DBLP:journals/corr/abs-2312-11805] to assign scores to each response following the established evaluation protocol, with ratings ranging from 1 to 10. For each test speech sample, the final evaluation score is computed as the ratio of its rating to that of a high-quality reference response. Formally, for $N$ evaluation instances (where $N=392$, corresponding to the size of our test set), the metric is defined as
$\text{Score}=\frac{1}{N}\sum_{i=1}^{N}\frac{s_{i}^{\text{test}}}{s_{i}^{\text{ref}}}.$
This ratio standardizes scores across different prompts and reduces variance caused by prompt differences, thereby improving stability and promoting better alignment with human judgments.
Models IA $C$ SF SN PC EA PeC KC Overall
English Evaluation Results
Qwen2.5-Omni 0.539 0.657 0.677 0.633 0.458 0.483 0.399 0.602 0.556
SpeechRole-Agent 0.806 0.957 0.881 0.791 0.608 0.689 0.608 0.883 0.778
Qwen2.5-Omni 0.531 0.608 0.700 0.666 0.489 0.518 0.417 0.567 0.562
SpeechRole-Agent 0.720 0.940 0.864 0.751 0.566 0.625 0.557 0.795 0.727
Chinese Evaluation Results
Qwen2.5-Omni 0.654 0.766 0.891 0.822 0.615 0.629 0.525 0.640 0.693
SpeechRole-Agent 0.908 0.935 0.896 0.849 0.763 0.800 0.739 0.859 0.843
Qwen2.5-Omni 0.599 0.773 0.884 0.700 0.578 0.580 0.450 0.707 0.659
SpeechRole-Agent 0.859 0.925 0.940 0.832 0.768 0.815 0.752 0.909 0.850
5 Experiments
This section presents a comprehensive experimental evaluation of SRPAs on SpeechRole-Eval. We evaluate a diverse set of representative baselines spanning both cascaded and end-to-end system paradigms, which reflect the two dominant design choices for current SRPAs. Experiments are conducted in both Chinese and English settings to assess multilingual performance and generalization. In addition, we analyze the effect of role-specific training by comparing a base end-to-end model with its counterpart further trained on SpeechRole-Data. Finally, we conduct a human–LLM agreement study to validate the reliability of the proposed automatic evaluation framework.
5.1 Evaluated Systems
We evaluate a total of seven baseline SRPAs, grouped into cascaded and end-to-end systems.
5.1.0.1 Cascaded Systems.
Cascaded SRPAs follow a core three-stage pipeline consisting of ASR, text-based reasoning, and TTS. Specifically, user speech is first transcribed using whisper-large-v3-turbo [DBLP:conf/icml/RadfordKXBMS23]. The transcribed text is then processed by a large language model for generating role-conditioned responses. We consider three open-source LLMs for this stage: Qwen3-8B [DBLP:journals/corr/abs-2505-09388], Llama-3.1-8B [DBLP:journals/corr/abs-2407-21783], and Mistral-7B-Instruct-v0.3 [DBLP:journals/corr/abs-2310-06825]. Finally, generated responses are converted into speech using F5-TTS [DBLP:journals/corr/abs-2410-06885]. The TTS component supports voice cloning based on reference audio, enabling the synthesized speech to match the target character’s timbre.
In addition to open-source cascaded systems, we include a proprietary cascaded baseline built upon the Alibaba Cloud API [alibabacloud], which integrates Paraformer Realtime ASR v2, Qwen-Plus-Character for role-conditioned text generation, and CosyVoice-v2 for speech synthesis. This system also supports reference-based voice cloning.
5.1.0.2 End-to-End Systems.
End-to-end SRPAs directly generate speech responses from speech inputs, avoiding explicit intermediate text representations. Such systems are expected to reduce error accumulation and latency, and are closer to the ideal formulation of SRPAs. We evaluate three end-to-end models: the proprietary GPT-4o Audio [DBLP:journals/corr/abs-2410-21276], and the open-source Qwen2.5-Omni-7B [DBLP:journals/corr/abs-2503-20215] and LLaMA-Omni [DBLP:conf/iclr/FangGZMZ025]. Notably, LLaMA-Omni only supports English and is therefore excluded from Chinese evaluations.
To further investigate the impact of role-specific speech data, we fine-tune Qwen2.5-Omni-7B on SpeechRole-Data, resulting in a specialized model referred to as SpeechRole-Agent. We compare its performance against the base model to assess the benefits of targeted SRPA training.
5.2 Overall Results and System Comparison
Table 2 summarizes the overall performance of all SRPAs on SpeechRole-Eval in English and Chinese. Overall, closed-source systems achieve the strongest results, with GPT-4o Audio consistently ranking at or near the top across languages. The cascaded system based on Alibaba Cloud API also performs competitively, particularly in Chinese. In contrast, open-source end-to-end models lag significantly behind across most metrics, resulting in substantially lower overall scores.
5.2.0.1 Interaction Ability.
IA and CC are strongly related to a model’s language understanding and reasoning ability, but our results show that they are also influenced by the speech modeling paradigm. Cascaded systems built on strong text-based LLMs, such as Qwen3-8B and Llama-3.1-8B, achieve consistently high scores on both metrics. However, end-to-end models may exhibit substantial degradation even when sharing the same text backbone. For example, LLaMA-Omni underperforms its cascaded Llama-3.1-8B counterpart on IA and CC, indicating that unified speech-text modeling can impair interaction quality if cross-modal integration is not sufficiently optimized. In contrast, GPT-4o Audio demonstrates that strong end-to-end SRPAs can still maintain robust interaction ability. Overall, these results suggest that IA and CC depend jointly on language modeling strength and effective speech-text integration.
5.2.0.2 Speech Quality and Expressiveness.
GPT-4o Audio demonstrates clear advantages in SF and SN, reflecting the benefits of tightly integrated speech generation. However, its gains in PC and EA are more limited, with several cascaded systems achieving comparable or even slightly stronger performance. This suggests that while end-to-end modeling improves basic perceptual speech quality, expressive control remains challenging. Overall, despite their theoretical advantages, current open-source end-to-end systems do not yet consistently surpass well-engineered cascaded approaches on expressive speech metrics.
5.2.0.3 Role-Playing Fidelity.
PeC and KC exhibit trends broadly aligned with interaction ability, with systems based on stronger language models generally maintaining more stable persona traits and role-consistent knowledge. However, similar to IA and CC, end-to-end models can show notable degradation even when built on comparable text backbones. For instance, LLaMA-Omni underperforms its cascaded counterpart in both PeC and KC, suggesting that unified speech-text modeling may weaken persona grounding and knowledge consistency when cross-modal alignment is insufficient. At the same time, GPT-4o Audio achieves competitive role fidelity, indicating that end-to-end designs can preserve character consistency when adequately optimized. Overall, role-playing fidelity depends on both the strength of the language model and the effectiveness of speech-text integration.
5.3 Effect of Role-Specific Training Data
To assess the impact of role-specific speech training, we compare the base Qwen2.5-Omni model with SpeechRole-Agent, which is fine-tuned on SpeechRole-Data covering 78 training roles. The evaluation is further divided into in-domain roles seen during training and out-of-domain roles that are held out, allowing us to examine both fitting and generalization behavior. The quantitative results are summarized in Table 3.
SpeechRole-Agent achieves substantial improvements over the base model in the in-domain setting across all metrics, with particularly large gains in IA, CC, PeC, and KC. This indicates that role-specific supervision effectively enhances structured role-playing behavior in end-to-end systems.
Importantly, these improvements largely persist in the out-of-domain setting. Although some degradation is observed compared to in-domain roles, SpeechRole-Agent consistently outperforms Qwen2.5-Omni across both languages and most evaluation dimensions, leading to significantly higher overall scores. This suggests that training on SpeechRole-Data improves not only role memorization but also general role-playing competence.
5.4 Human–LLM Agreement Analysis
To assess whether the proposed LLM-based evaluation aligns with human judgment, we conduct a human–LLM agreement study on a subset of SpeechRole-Eval. We focus on Chinese role-playing dialogues to facilitate consistent expert assessment under a shared linguistic context. Half of the Chinese evaluation instances are uniformly sampled, covering all dialogue types and role categories, and all baseline systems are evaluated on each instance.
Three expert annotators independently assign absolute scores on a 1–10 scale following the evaluation dimensions defined in Section 4.1. Annotators are blinded to system identities and LLM scores. Human ratings are aggregated by averaging scores across annotators and evaluation instances for each baseline system.
We evaluate agreement by computing Spearman’s rank correlation between the aggregated human scores and the automatic scores produced by SpeechRole-Eval at the system level. Strong correlations are observed across evaluation dimensions, indicating that the proposed LLM-based evaluation produces rankings aligned with human judgment and can serve as a proxy for expert evaluation. Detailed annotation protocols and additional analysis are provided in Appendix F.
6 Conclusion
This work introduces SpeechRole, a unified framework for advancing SRPAs through both large-scale data and systematic evaluation. We present SpeechRole-Data, a speech-to-speech role-playing dataset covering 98 roles with rich timbral and prosodic variation, and SpeechRole-Eval, a multidimensional benchmark that directly evaluates generated speech across interaction ability, speech expressiveness, and role-playing fidelity.
Through extensive experiments on both cascaded and end-to-end systems, we provide a comprehensive analysis of current SRPAs. Our results show that while state-of-the-art end-to-end models achieve strong fluency and naturalness, challenges remain in prosody control and emotional expressiveness. Moreover, interaction ability and role-playing fidelity are largely constrained by the underlying text-based language models, highlighting a key bottleneck for future progress.
We further demonstrate that the proposed LLM-based evaluation framework aligns well with expert human judgments, supporting its use as a scalable and reliable alternative to human evaluation. We hope SpeechRole will serve as a foundation for future research on expressive, controllable, and character-consistent speech role-playing systems.
7 Limitations
Despite its contributions, this work has several limitations. First, although SpeechRole-Data covers a diverse set of roles and speaking styles, the number of languages remains limited, with experiments primarily focusing on English and Chinese. Extending the dataset and evaluation framework to more languages and cultural contexts is an important direction for future work.
Second, our automatic evaluation relies on a strong proprietary LLM as the judge. While human correlation analysis shows high agreement, the evaluation may still inherit biases or blind spots from the underlying model. Developing fully open-source judges or complementary objective speech metrics remains an open challenge.
Third, the current evaluation focuses on system-level performance and averaged behavior. It does not explicitly assess fine-grained controllability, long-term character drift, or robustness under adversarial or ambiguous role instructions. These aspects are critical for real-world deployment of SRPAs and warrant further investigation.
Finally, while SpeechRole-Agent demonstrates clear gains through task-specific training, it is built upon an existing end-to-end model. Exploring alternative architectures, training objectives, and tighter integration between language and speech modeling may further improve expressive and faithful role-playing in future systems.
8 Ethics Statement
This work does not involve the release of newly annotated human-labeled datasets. Human annotation is only conducted for evaluation purposes in the Human–LLM agreement study, and the resulting annotations are not included in the released resources. All human evaluations are performed by expert annotators selected from the research team, following a predefined evaluation protocol. No personal data is collected, and all evaluated content is derived from fictional role-playing dialogues. Therefore, we believe this work raises no significant ethical concerns.
Appendix A AI Assistants in Research or Writing
In preparing this manuscript, AI assistants were employed solely to assist with refining the clarity, style, and readability of certain text segments. They were not involved in designing the study, developing or implementing the methodology, collecting or analyzing data, or generating the primary scientific contributions. All substantive research decisions, analyses, and conclusions are fully the responsibility of the authors.
Appendix B Evaluation Metrics Details